### v2

@@
[orcid=$0$-0001-9718-4464]
[orcid=$0$-0001-7975-3985]
@@
Ethan Kane Waters ethan.waters@jcu.edu.au [ [    Carla Chia-Ming Chen carla.ewels@jcu.edu.au    Mostafa Rahimi Azghadi mostafa.rahimiazghadi@jcu.edu.au
@@ 1 Introduction
Disease presents a formidable hurdle to maximising yield in the sugarcane industry with Ratoon stunting disease (RSD) emerging as a key contributor globally (Chakraborty et al., 2024). The bacterium Leifsonia xyli subsp. xyli (Lxx) is the primary causal agent of RSD, which is primarily propagated with contaminated cutting implements (Davis et al., 1984). RSD infection can result in substantial yield losses (up to 60%), depending on sugarcane variety and water availability during periods of growth (Bailey and Bechet, 1986; Chakraborty et al., 2024; Davis and Bailey, 2000; Croft et al., 2000). This results in pronounced economic repercussions, with an estimated annual loss of \$25 million observed across 87,000 hectares monitored in Australia in 2019(Magarey et al., 2021). The lack of external symptoms (Chakraborty et al., 2024; Davis and Bailey, 2000; Croft et al., 2000; Sugar Research Australia, 2021) has made the detection and the management of RSD a significant challenge. It relies heavily on laboratory diagnostic techniques such as PCR, qPCR, LAMP, or LSB-qPCR for RSD detection (Carvalho et al., 2016; Fegan et al., 1998; Ghai et al., 2014; Young et al., 2016). Collecting the field samples required for these techniques is both time-consuming and costly, especially for large-area detection. This underscores the need to develop a more efficient method for large-scale detection of RSD.
@@
Freely available satellite-based remote sensing offers a cost-effective and efficient alternative to the traditional, resource-intensive methods of identifying and managing disease in sugarcane (Waters et al., 2024). In particular, free publicly accessible multispectral satellite data can alleviate the financial burden of purchasing expensive spectrometers images, and promote the widespread adoption of this advanced technology.
@@
The aim of this research is therefore to investigate the efficacy of various machine learning algorithms in classifying asymptomatic RSD across several sugarcane varieties using freely available satellite data, including identifying which vegetation indices are important for the diagnosis of RSD if at all. To date, this study represents the first effort to classify an asymptomatic disease in sugarcane through satellite-based spectroscopy (Waters et al., 2024). It is also the first instance of classifying RSD with machine learning and spectroscopy (Waters et al., 2024).
@@ 2.1 Overall Process Summary
The project comprises three stages: ground-truth data collection, data preprocessing, and ML development. In stage 1, RSD data is collected from sugarcane blocks. For stage 2, sentinel-2 images with pre-applied atmospheric correction undergo preprocessing before pixelwise values are extracted and labeled from raw spectral bands and vegetation indices. In the third stage, the dataset is divided into an 80:20 train-test split, with 10-fold validation applied to the training set for hyperparameter tuning. The top-performing models are subsequently evaluated on the test set by performing bootstrapping with 5000 samples. Permutation testing is then performed to ensure model performance is significantly better than a null distribution. The steps are detailed in the following sections.
@@ 2.2 Study Location & Ground Truth Data Collection
Sampling occurred between February and March 2022 across 72 sugarcane blocks situated in the Herbert region of Queensland, Australia. The acquisition of a ground-truth dataset detailing the disease status and variety of sugarcane was conducted by trained field agronomists at Herbert Cane Productivity Services Limited (HCPSL).
Two types of sugarcane blocks were sampled; farmer-owned blocks and HCPSL-owned seed production blocks. HCPSL-owned seed production blocks exist to provide farmers the opportunity to purchase disease-free sugarcane seeds due to the strict hygiene and disease testing requirements. HCPSL followed their standard RSD sampling protocol for both types of blocks. In seed production plots samples were collected every 20m in a grid pattern. In Farmer-owned blocks, 12 samples were collected from the four-corners of the block, and four from randomly selected locations within the block. The juice was then extracted from the samples and underwent qPCR testing at Sugar Research Australia (SRA) laboratories. The block is classified as RSD Positive if any sample within that block returned as RSD positive. To avoid potential false negatives in farmer-owned blocks due to sparse sampling, only HCPSL seed production blocks were used for RSD-negative samples.
The shapefile corresponding to the sample blocks and the corresponding sugarcane variety was provided by HCPSL. In total, data were collected from 72 blocks, which comprised five sugarcane varieties, Q200, Q208, Q240, Q253 and SRA14.
@@ 2.3 Multispectral Satellite Image Acquisition and Data Labelling
The multispectral data used in the study were captured by Sentinel-2 on 27th February 2022, as it had the least cloud coverage during the sampling period. Since RSD is transmitted through contaminated tools during planting or harvesting, and the study period was during the plant growth period, we assumed that the infection status remained unchanged during the study period, hence only one day of image data was used in this study.
The Sentinel-2 data processing pipelines used the QGIS Python API to transform the coordinate reference system (CRS) of the Sentinel-2 products to match the CRS of the ground truth data shapefile representing the farms of interest. All Sentinel-2 pixels within the geometry of the 72 sampled blocks described in the shapefile were labelled with the disease status and variety of the block. Each labelled pixel within the shapefile geometries was utilised as a separate observation. Table 2 lists the different sugarcane varieties, with their respective RSD positive/negative pixel numbers.
@@
2.4 Vegetation Indices
The causal bacteria of RSD infect the xylem vessels responsible for water transport in sugarcane, reducing water absorption and resulting in vegetation stress (Davis et al., 1984; Chakraborty et al., 2024). Previous research has demonstrated that Sentinel-2’s bands 11 and 12 in the Short-wave Infrared Region (SWIR) are sensitive to moisture absorption in vegetation, making them potential predictors for RSD (Fensholt and Sandholt, 2003; Liu et al., 2021; Wang et al., 2017). Additionally, Near Infrared Region (NIR) bands and vegetation indices derived from them have previously been utilised as indicators of vegetation health or stress (Apan et al., 2004, 2003; Narmilan et al., 2022). An independent t-test with a significance value of 0.05 was performed for each variety and spectral band, adjusting for type I error with Bonferroni’s Correction, to provide an indication of which bands and vegetation indices may be effective predictors of RSD.
Although the statistically significant bands varied between varieties, bands B08A, and B12 were consistent across a majority of varieties. This supports the notion that vegetation moisture and stress are key indicators for diagnosing RSD. Accordingly, the vegetation indices selected for this study focus on vegetation moisture and stress with several general health indicators, as shown in Table 3, were chosen to emphasise these features. Specifically, the DWSIs are of interest given their demonstrated effectiveness for detecting sugarcane disease with satellite-based remote sensing Apan et al. (2003); Dutia et al. (2006) by focusing on vegetation moisture. The original indices were formulated with hyperspectral satellite imagery, specifically targeting R1600 to measure water absorption. In this study, these indices were replicated as closely as possible with Sentinel-2 data. For example, Sentinel-2’s band 11 is approximately centered at R1600, was selected for its ability to capture similar moisture-related signals. Given that bands 11 and 12 are both sensitive to moisture (Fensholt and Sandholt, 2003; Liu et al., 2021; Wang et al., 2017), and the statistical significance observed in the analysis for band 12, the original DWSIs were adapted to include band 12. This led to the creation of several new indices — DWSI 6, 7, and 8, — that integrate both bands to enhance their sensitivity to RSD symptoms. To our knowledge, these modified indices have not been previously utilised. Additionally, Sentinel-2’s band 5, a red-edge band known to improve vegetation classification accuracy (Kussul et al., 2017; Qiu et al., 2017), was incorporated into DWSI-8 in an attempt to further improve its predictive performance.
The spectral bands of the satellite image captured on the 27th February 2022 were utilised to compute the 19 vegetation indices listed in Table 3 with the QGIS Python API for each pixel in the dataset.
@@ 2.4 Vegetation Indices
This Study Disease-Water Stress Index 6 (DWSI-6) $\frac{12+11}{8A}$
This Study Disease-Water Stress Index 7 (DWSI-7) $\frac{12+11}{4}$
This Study Disease-Water Stress Index 8 (DWSI-8) $\frac{12+11}{4+5}$
@@ 2.5 Machine Learning Algorithm Development
In this study, we compared and contrasted five machine learning algorithms to predict if a pixel is RSD positive. These are Random Forest (RF), Logistic Regression (LR), Quadratic Discriminant Analysis (QDA), Gradient Boosting (GB) and Radial Basis Function Support Vector Machine (SVM-RBF). Given that the sugarcane variety can be potentially unknown to the users, We developed two sets of models for the scenarios when the sugarcane variety is known and unknown. The final dataset seen in Table 2 is largely unbalance, therefore, the dataset was down-sampled to achieve equal class distribution, to mitigate potential bias.
@@
The dataset was partitioned into 80% for training and 20% for testing. Standard scaling was applied to the entire dataset, with the scaling parameters fitted exclusively on the training set. Hyperparameters were tuned using 10-fold cross-Validation with "Halving Grid Search" on the 80% training set. HalvingGridSearchCV is a resource-efficient version of GridSearchCV from the scikit-learn library for hyperparameter tuning. It uses successive halving, which starts with many candidate parameter combinations and progressively eliminates the least promising ones, allocating more resources as the number of iterations increases (Scikit-learn Developers, 2024).
The model performance was evaluated by bootstrapping the unseen test set with 5000 bootstrap samples, yielding distributions for accuracy, precision, and recall. After the model evaluation, each machine learning algorithm was trained with 10-fold cross-validation on the entire dataset, to find the most appropriate hyperparameters for the entire dataset, after the performance has been determined.
@@
To assess the significance of the ML algorithms’ ability to detect RSD, a permutation test was conducted. The pixel labels were shuffled, and the models were retrained using the previously determined hyperparameters. This process was repeated 1000 times to generate a null distribution of accuracy, precision, and recall, reflecting model performance when there is no true association between the features and disease status. By comparing the original performance distribution to the null distribution, the statistical significance of the observed model performance can be evaluated.
2.6 Software and Libraries

### v3

@@
3]organization=Department Of Primary Industries, addressline=203 Tor Street, city=Toowoomba, postcode=4350, state=QLD, country=Australia
[orcid=0000-0001-9718-4464]
[orcid=0000-0001-7975-3985]
@@
Ethan Kane Waters ethan.waters@dpi.qld.gov.au [ [ [    Carla Chia-Ming Chen carla.ewels@jcu.edu.au    Mostafa Rahimi Azghadi mostafa.rahimiazghadi@jcu.edu.au
@@ 1 Introduction
Disease presents a formidable hurdle to maximising yield in the sugarcane industry with Ratoon stunting disease (RSD) emerging as a key contributor globally (Chakraborty et al., 2024). The bacterium Leifsonia xyli subsp. xyli (Lxx) is the primary causal agent of RSD, which is primarily propagated with contaminated cutting implements (Davis et al., 1984). RSD infection can result in substantial yield losses (up to 60%), depending on sugarcane variety and water availability during growth periods (Bailey and Bechet, 1986; Chakraborty et al., 2024; Davis and Bailey, 2000; Croft et al., 2000). This results in pronounced economic repercussions, with an estimated annual loss of \$25 million observed across 87,000 hectares monitored in Australia in 2019(Magarey et al., 2021). The lack of external symptoms (Chakraborty et al., 2024; Davis and Bailey, 2000; Croft et al., 2000; Sugar Research Australia, 2021) has made the detection and the management of RSD a significant challenge. It relies heavily on laboratory diagnostic techniques such as Polymerase Chain Reaction (PCR), Quantitative Polymerase Chain Reaction (qPCR), Loop-Mediated Isothermal Amplification (LAMP), or Leaf Sheet Biopsy Quantitative Polymerase Chain Reaction (LSB-qPCR) for RSD detection (Carvalho et al., 2016; Fegan et al., 1998; Ghai et al., 2014; Young et al., 2016). Collecting the field samples required for these techniques is both time-consuming and costly, especially for large-area detection. This underscores the need to develop a more efficient method for large-scale detection of RSD.
@@
Freely available satellite-based remote sensing offers a cost-effective and efficient alternative to the traditional, resource-intensive methods of identifying and managing disease in sugarcane (Waters et al., 2025). In particular, free publicly accessible multispectral satellite data can alleviate the financial burden of purchasing expensive spectrometers images, and promote the widespread adoption of this advanced technology.
@@
The primary objective of this research is to evaluate the effectiveness of various machine learning algorithms in classifying asymptomatic RSD in multiple varieties of sugarcane with freely available multispectral satellite data. Additionally, the study seeks to identify which vegetation indices are most important for diagnosing RSD. To the best of our knowledge, this study represents the first effort to classify an asymptomatic disease in sugarcane with satellite-based spectroscopy and is the first instance of diagnosing RSD with machine learning and spectroscopy (Waters et al., 2025).
@@ 2.1 Overall Process Summary
The project comprises three stages: ground-truth data collection, data preprocessing, and ML development. In stage one, RSD data is collected from sugarcane blocks. For stage two, sentinel-2 images with pre-applied atmospheric correction undergo preprocessing before pixelwise values are extracted and labeled from raw spectral bands and vegetation indices. In the third stage, the dataset is divided into an 80:20 train-test split, with 10-fold validation applied to the training set for hyperparameter tuning. The top-performing models are subsequently evaluated on the test set by performing bootstrapping with 5000 samples. Permutation testing is then performed to ensure model performance is significantly better than a null distribution. The steps are detailed in the following sections.
@@ 2.2 Study Location & Ground Truth Data Collection
Sampling to develop the ground truth dataset for this study was conducted by trained field agronomists at Herbert Cane Productivity Services Limited (HCPSL) between February and March 2022 across 76 sugarcane blocks in the Herbert region of Queensland, Australia. The dataset details the RSD status and variety of the sugarcane in each of the 76 blocks, which are comprised of five sugarcane varieties, Q200, Q208, Q240, Q253 and SRA14.
Two types of sugarcane blocks were sampled; farmer-owned blocks and HCPSL-owned seed production blocks. HCPSL-owned seed production blocks exist to provide farmers the opportunity to purchase disease-free sugarcane seeds due to strict hygiene and disease testing requirements. HCPSL followed their standard RSD sampling protocol for both types of blocks. In seed production plots samples from a sugarcane stick was collected every 20m in a grid pattern. In Farmer-owned blocks, samples from twelve sugarcane sticks were collected from the four-corners of the block, and four from randomly selected locations within the block. The xylem extraction method was employed, in which juice from each sugarcane stick was extracted into a vial using pressure generated by an air compressor. The extracted juice from all samples was then frozen and subsequently analysed with qPCR testing at Sugar Research Australia (SRA) laboratories. The entire block is classified as RSD Positive if any sample within that block returned as RSD positive, and only labelled RSD Negative if all samples return a negative result. A visualisation of the field sampling method for the two block types is shown in Figure 1.
To establish reliable ground truth data for model training and validation, RSD negative samples were sourced exclusively from HCPSL seed cane blocks that returned negative qPCR test results, while RSD positive samples were sourced from grower blocks that returned positive qPCR test results. The more intensive sampling protocols implemented by HCPSL staff to meet strict disease control requirements led to different sampling densities between the two block types. However, this does not introduce bias into the dataset. The study labels data at a block level and consequently does not analyse within block variation which could be influenced by different sampling densities. Additionally, RSD Negative samples were only drawn from HCPSL seed blocks with the higher sampling density to increase confidence that these blocks were truly absent of RSD. This conservative approach to labelling negative samples is appropriate given the asymptomatic and contagious nature of RSD, where undetected infections pose a risk of mislabelling. In contrast, RSD positive samples were drawn only from grower blocks where qPCR testing confirmed disease presence. While these blocks were sampled less intensively, this does not compromise the validity of the positive labels, as only confirmed cases were included. It is possible that some infected blocks were missed due to lower sampling density and therefore excluded from the dataset entirely. Overall, the differing sampling strategies were intentionally structured to minimise label noise and enhance the reliability of both positive and negative class assignments.
A shapefile corresponding to the sampled blocks with the sugarcane variety and RSD status determined in the laboratory testing was produced by HCPSL GIS officer Rod Nielson. However, due to a nondisclosure agreement with the farmers, the shapefile and precise maps of the sampled areas cannot be published.
Figure 1: Visual representation of the field sampling for both block types where the red marker indicates a sampling location.
@@ 2.3 Multispectral Satellite Image Acquisition and Data Labelling
The multispectral data used in the study were captured by Sentinel-2 on 27th February 2022, as it had the least cloud coverage during the sampling period. Since RSD is transmitted through contaminated tools during planting or harvesting, and the study period was during the plant growth period, we assumed that the infection status remained unchanged during the study period. Only one day of image data was used in this study.
The Sentinel-2 data processing pipeline used the QGIS Python API to transform the coordinate reference system (CRS) of the Sentinel-2 products to match the CRS of the ground-truth shapefile representing the farms of interest. The ground-truth data was then used to extract the 76 sampled blocks from the Sentinel-2 imagery, ensuring that each pixel within the block geometries was labelled with both disease status and variety. Figure 2 shows an example block across Sentinel-2 bands, highlighting within-field variation captured in the multispectral data. Each labelled pixel within the shapefile geometries was subsequently utilised as a separate observation. Table 2 summarises the number of ground-truth pixel observations at 20m resolution for each sugarcane variety.
((a)) B02 (490 nm)
((b)) B03 (560 nm)
((c)) B04 (665 nm)
((d)) B05 (705 nm)
((e)) B06 (740 nm)
((f)) B07 (783 nm)
((g)) B8A (865 nm)
((h)) B11 (1610 nm)
((i)) B12 (2190 nm)
Figure 2: Sentinel-2 multispectral bands for an example block, with each panel showing a different band. Pixel values have been normalised to highlight within-field variation.
@@
The distributions of reflectance values across Sentinel-2 bands vary with both disease status and variety. Figure 3 illustrates the spectral response of each variety. For example, in variety Q240, RSD-positive pixels generally exhibit lower reflectance in the red-edge and near-infrared regions (bands 6–8) compared with RSD-negative pixels, whereas varieties such as SRA14 and Q253 show higher median reflectance for RSD-positive pixels.
Figure 3: Reflectance distributions across Sentinel-2 bands for five sugarcane varieties and an overall panel. Outliers are omitted for clarity. Healthy and diseased crops are colour-coded (green = healthy, orange = diseased).
@@
This Study Disease-Water Stress Index 6 (DWSI-6) $\frac{B12+B11}{B8A}$
This Study Disease-Water Stress Index 7 (DWSI-7) $\frac{B12+B11}{B4}$
This Study Disease-Water Stress Index 8 (DWSI-8) $\frac{B12+B11}{B4+B5}$
2.4 Vegetation Indices
The causal bacteria of RSD infects the xylem vessels responsible for water transport in sugarcane, reducing water absorption and resulting in vegetation stress (Davis et al., 1984; Chakraborty et al., 2024). To detect these internal symptoms, vegetation indices were selected that measure key indicators of plant health, as shown in Table 3. Several indices were chosen to assess vegetation moisture content (Gao, 1996; Apan et al., 2004), with the intention highlighting reduced water uptake caused by RSD infection of the xylem. Others focus on assessing symptoms of plant stress (Merzlyak et al., 1999; Apan et al., 2003; Rouse Jr et al., 1973), indicated by changes in photosynthetic activity or chlorophyll levels. In addition to vegetation indices, previous research has demonstrated that the raw Sentinel-2’s bands in the Short-wave Infrared Region (SWIR) are sensitive to moisture absorption in vegetation (Fensholt and Sandholt, 2003; Liu et al., 2021; Wang et al., 2017), while bands in the Near-Infrared (NIR) region have been previously used as indicators of vegetation stress (Segarra et al., 2020), making both potential predictors for RSD. Together, the indices and bands provide a comprehensive net to capture the expected physiological impacts of RSD on sugarcane.
To optimise class separation, strategic modifications of vegetation indices were explored. The alteration of existing indices was guided by first identifying statistically significant bands for diagnosing RSD with an independent t-test (significance level = 0.05) for each variety and spectral band, with Bonferroni’s correction applied to adjust for type I error. Although statistically significant bands varied between varieties, bands B08A and B12 were consistent across most varieties, supporting the connection between vegetation moisture and stress as key indicators for diagnosing RSD. With the identification of statistically significant bands, opportunities arise to modify key vegetation indices to enhance class separation. Among the most relevant indices are the Disease Water Stress Indices (DWSIs), which have proven effective in detecting sugarcane disease through satellite-based remote sensing Apan et al. (2003). The DWSIs were designed to target water absorption at R1660 (Sentinel-2’s B11). Incorporating both B11 and B12, given their vegetation moisture absorption sensitivity (Fensholt and Sandholt, 2003; Liu et al., 2021; Wang et al., 2017) and B12’s statistical significance, led to modified versions of these indices for Sentinel-2, DWSI 6 and 7. Additionally, DWSI 7 was further modified to include the red-edge band, B5, as the red-edge region is sensitive to stress-related changes in vegetation and has previously been effective for vegetation classification (Kussul et al., 2017; Qiu et al., 2017). This modification resulted in the creation of DWSI 8, designed to better capture stress-related signals in vegetation as a result of RSD.
The spectral bands of the satellite image captured on the 27th February 2022 were utilised to compute the 19 vegetation indices listed in Table 3 with the QGIS Python API for each pixel in the dataset. A visualisation of several key vegetation indices for an example block is shown in Figure LABEL:fig:vi_block, illustrating patterns relevant to sugarcane disease detection.
((a)) DWSI-1
((b)) DWSI-2
((c)) DWSI-3
((d)) DWSI-4
((e)) DWSI-5
((f)) DWSI-6
((g)) DWSI-7
((h)) NDMI
((i)) NDVI
Figure 4: Nine vegetation indices for a sample block that are historically significant for vegetation monitoring or were found to be important for detecting RSD. Pixel values have been normalised to highlight within-field variation.
@@ 2.5 Machine Learning Algorithm Development
In this study, we compared five machine learning algorithms to predict the RSD status of a pixel. These are Random Forest (RF), Logistic Regression (LR), Quadratic Discriminant Analysis (QDA), Gradient Boosting (GB) and Radial Basis Function Support Vector Machine (SVM-RBF). Tree-based methods, such as RF and GB, were included due to their previous success in similar remote sensing studies. These models are particularly effective for handling high-dimensional data, non-linearity, and interactions among features, making them well-suited for RSD detection. SVM-RBF kernel was included due to its ability to effectively model complex non-linear relationships, which are often present in multispectral data. LR was included due to its simplicity and interpretability. LR is useful for assessing whether the problem could be adequately solved with linear methods without resorting to more complex models. QDA was tested as it is capable of modelling quadratic decision boundaries, which can be useful for capturing more complex class separations that may exist in the data.
@@
Given that the sugarcane variety can be potentially unknown to the users, two sets of models were developed, one for when the sugarcane variety is known and the other when its unknown. The final dataset seen in Table 2 is largely unbalance, therefore, the dataset was down-sampled to achieve equal class distribution, to mitigate potential bias.
Figure 5: Visual guide to the machine learning workflow. The dataset was split into training and testing sets. The training set underwent 10-fold cross-validation to select the best model. Model performance was then assessed by bootstrapping the test set 5,000 times, followed by permutation testing.
@@
The dataset was partitioned into 80% for training and 20% for testing. Standard scaling was applied to the entire dataset, with the scaling parameters fitted exclusively on the training set. Hyperparameters were tuned using 10-fold cross-Validation with "Halving Grid Search" on the 80% training set. HalvingGridSearchCV is a resource-efficient version of GridSearchCV from the scikit-learn library for hyperparameter tuning. HalvingGridSearchCV is an implementation of the successive halving algorithm, introduced by Jamieson and Talwalkar (2016), which starts with many candidate parameter combinations and progressively eliminates the least promising ones, allocating more resources as the number of iterations increases (Scikit-learn Developers, 2024). Table 4 provides a summary of the key parameters adjusted, including the search space explored.
The model’s performance was evaluated by bootstrapping the unseen test set with 5000 bootstrap samples, generating distributions for the chosen performance metrics, accuracy, precision, and recall (Ting, 2010). Accuracy provides an overall measure of correct predictions, while precision and recall offer insights into the model’s ability to correctly identify positive cases and avoid false positives, respectively. In the context of large-scale disease detection, evaluating false positives and false negatives is critical to ensuring an appropriate balance is achieved. High false positive rates could lead to unnecessary interventions by agronomists, wasting time, diverting resources, and reducing the economic benefit of the system. After model evaluation, each machine learning algorithm was trained with 10-fold cross-validation on the entire dataset, to find the most appropriate hyperparameters for the entire dataset, after the performance has been determined. A visualisation of this machine learning workflow is shown in Figure 5.
Table 4: Summary of hyperparameters search space utilised in halving grid search.
Model Parameter Search Space
SVM C 0.1, 1, 10, 100, 1000, 10000, 100000, 1000000
gamma 0.001, 0.01, 0.1, 1, 10, 100
RF n_estimators 50, 100, 200, 500, 1000, 1500
max_depth None, 10, 15, 20, 30
GB n_estimators 50, 100, 200, 500, 1000, 1500
learning_rate 0.01, 0.1, 0.2, 0.3
max_depth 3, 5, 7, 10
LR penalty ’l1’, ’l2’
solver ’liblinear’, ’saga’
gamma 0.001, 0.01, 0.1, 1, 10, 100
QDA N/A N/A
The Gini importance of the final RF and GB models was extracted to identify the most important vegetation indices and Sentinel-2 multispectral bands for RSD detection. Additionally, to indicate important features for RSD detection, independent t-tests (significance level = 0.05) were performed for each vegetation index, applying Bonferroni’s correction to adjust for type I error. Although the results are included in the appendix, this approach does not account for correlations among the indices or interactions between them. Therefore, the findings should be interpreted with caution.
@@
To assess the significance of the ML algorithms’ ability to detect RSD, a permutation test was conducted. The pixel labels were shuffled, and the models were retrained using the previously determined hyperparameters. This process was repeated 1000 times to generate a null distribution of accuracy, reflecting model performance when there is no true association between the features and disease status. By comparing the original performance distribution to the null distribution, the statistical significance of the observed model performance can be evaluated.
2.6 Software, Libraries & Experimental Environment